OpenStudy (anonymous):

Simplify √-48 a. -4√3 b. 4√-3 c. 4i√3 d. 4√3i

OpenStudy (anonymous):

\[i4\sqrt{3}\]

OpenStudy (anonymous):

d

OpenStudy (anonymous):

cause \[\sqrt{-1}=i\]

OpenStudy (petiteme):

Its D. :)

OpenStudy (anonymous):

\[\sqrt{48}i\]